At what temperature will a solution that is composed of 0.73 moles of glucose (C6H12O6) in 650mL of water (H2O) begin to boil?
1
Expert's answer
2021-10-25T03:49:11-0400

Kb for water = 0.52 K kg mol -1

boiling points of pure water = 373.15k

650/18.015= 36.1moles

= 0.52×0.73×36.1= 13.7

13.7+373.15= 386.85K